# Graphon Quantum Filtering systems

### v1

@@ Abstract
We consider a non-exchangeable system of interacting quantum particles with mean-field type interactions, subject to continuous measurement on a class of dense graphs. In the mean-field limit, we derive a graphon-based quantum filtering system of equations, study its well-posedness, and establish a propagation of chaos result for multiple bosonic systems with blockwise interactions. We then discuss applications to quantum graphon games and quantum state preparation.
81S25, 81S22, 60K35, 60H30, 49N80,
quantum interacting-particle systems,
quantum filtering,
mean-field graph limits,
keywords:
[class=MSC]
keywords:
[…]
, , and
@@ 1 Introduction
The subject matter of this paper, is to introduce a formal framework for systems of quantum particles subjected to continuous measurement and their mean-field graph limits. The primary motivation for such models lies in their capacity to rigorously formulate dynamic quantum games and, by extension, quantum mean-field games (QMFGs) [42, 43, 40].
In fact in standard closed quantum mechanics the dichotomy between the state evolution (Schrödinger equation) and the measurement postulate (Born rule) prevents the observation of significant dynamics, this is known as the Zeno effect [53]. To overcome this, one must consider open quantum systems subjected to indirect measurement, where the natural way to incorporate continuous measurements is provided by the quantum filtering framework [13, 19, 20]. The a posteriori state evolution is then described by a modified Schrödinger equation of stochastic nature, known as the Belavkin equation or the quantum filtering equation. Depending on the nature of the measurement detection, two types of equations can be derived:
@@
    $\mathrm{d}\rho_{t}$ $=-\mathrm{i}[\tilde{H},\rho_{t}]\mathrm{d}t+\left(L\rho_{t}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\rho_{t}\}\right)\mathrm{d}t+\sqrt{\eta}\left(L\rho_{t}+\rho_{t}L^{\dagger}-\mathrm{t}\mathrm{r}\big{(}(L+L^{\dagger})\rho_{t}\big{)}\rho_{t}\right)\mathrm{d}W_{t},$
    where $(W_{t})_{t\geq 0}$ is a Wiener process, with the observation process
@@
    where $({N}_{t})_{t\geq 0}$ is a counting process with stochastic intensity depending on the state given by $\int_{0}^{t}\mathrm{t}\mathrm{r}(L\rho_{s}L^{\dagger})\mathrm{d}s$, with the observation process $Y_{t}=N_{t}.$
In this paper we focus only on the diffusive case, we will specify the significance of the terms later.
Initially, these equations were derived using the theory of quantum stochastic calculus [56]. However, they can also be obtained within the framework of standard quantum mechanics as an appropriate scaling limit in a repeated measurement setup (see [58, 57, 59]), or via the theory of instruments (see [5, 7]).
In his seminal series of papers, Kolokoltsov [41, 43, 40] developed the quantum mean-field games using the $N$-body quantum filtering equation as a starting point:
$\mathrm{d}\boldsymbol{\rho}_{t}^{N}$ $=-\mathrm{i}{\Big{[}\sum_{l=1}\mathbf{\tilde{H}}_{l}+\frac{1}{N}\sum_{l=1}\sum_{l^{\prime}>l}\mathbf{A}_{ll^{\prime}},\boldsymbol{\rho}_{t}^{N}\Big{]}}\mathrm{d}t+\sum_{l=1}^{N}\Big{(}{\bf L}_{l}\boldsymbol{\rho}_{t}^{N}{\bf L}^{\dagger}_{l}-\frac{1}{2}\big{\{}{\bf L}^{\dagger}_{l}{\bf L}_{l},\boldsymbol{\rho}_{t}^{N}\big{\}}\Big{)}\mathrm{d}t$
@@
where $\boldsymbol{\rho}^{N}_{t}$ represents the state of the $N$-body system which lives in huge tensorized space. Unlike in the classical framework, due to entanglement between quantum particles, we cannot separate the individual dynamics of the particles, and the concept of an empirical measure becomes meaningless in the quantum setting. Therefore, one must adopt a different methodology to establish propagation of chaos results.
Kolokoltsov adapted the Pickl method [60] to the stochastic setting to derive the corresponding limiting equation,
$\mathrm{d}\gamma_{t}$ $=-\mathrm{i}\Bigl{[}\tilde{H}+A^{\mathbb{E}[\gamma_{t}]},\gamma_{t}\Bigr{]}\mathrm{d}t+\Bigl{(}L\gamma_{t}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\gamma_{t}\}\Bigr{)}\mathrm{d}t+\Bigl{(}L\gamma_{t}+\gamma_{t}L^{\dagger}-\mathrm{tr}\bigl{(}(L+L^{\dagger})\gamma_{t}\bigr{)}\gamma_{t}\Bigr{)}\mathrm{d}W_{t}.$
@@
This leads to a new McKean–Vlasov equation formulated on an infinite-dimensional complex-valued Hilbert space.
Primary contributions
The aim of this paper is twofold: First, we extend the formalism of observed interacting quantum particles to the heterogeneous case. By exploiting the theory of graphons, we derive the following graphon-based quantum filtering system:
@@
$\mathrm{d}\gamma_{t,u}$ $=-\mathrm{i}\Bigl{[}\tilde{H}+\int_{0}^{1}w(u,v)A^{\mathbb{E}[\gamma_{t,v}]}\mathrm{d}v,\gamma_{t,u}\Bigr{]}\mathrm{d}t+\Bigl{(}L\gamma_{t,u}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\gamma_{t,u}\}\Bigr{)}\mathrm{d}t$
$\quad+\sqrt{\eta}\Bigl{(}L\gamma_{t,u}+\gamma_{t,u}L^{\dagger}-\mathrm{tr}\bigl{(}(L+L^{\dagger})\gamma_{t,u}\bigr{)}\gamma_{t,u}\Bigr{)}\mathrm{d}W^{u}_{t},\quad u\in I=[0,1],$
where $\mathbf{W}=\{W^{u}:u\in[0,1]\}$ is a collection of pairwise independent Wiener processes. To address the question of the well-posedness of the resulting graphon quantum filtering system, we specify a probability space that ensures the joint measurability of $\mathbf{W}$ via a Fubini extension [66]. We consider each $\gamma_{t,u}$ as taking values in the Liouville space $\mathcal{L}_{2}(\mathbb{H})$, which forms a Hilbert space [49]. This allows us to formulate the system as an equation in an appropriate Hilbert space, following the approach in [28].
To apply the standard fixed-point argument for the McKean–Vlasov equation, we parametrize the law dependence by a deterministic function and construct a weak solution via a Girsanov transformation, starting from the strong solution of the linearized equation. To ensure positivity, we adapt the infinite-dimensional Hilbert space argument from [51]. We then prove pathwise uniqueness and invoke the Hilbert-space formulation of the Yamada–Watanabe theorem to obtain a strong solution. Finally, to complete the well-posedness result, we establish stability with respect to the graphon function.
@@
Second, to the best of our knowledge, the mean-field limit for non-exchangeable quantum systems has not yet been studied. In this work, we focus on the special case in which the graphon limit is block-wise and the measurement efficiency is perfect $(\eta=1)$. In this setting, the limiting system can be described by a single representative state for each block. We extend Pickl’s method by introducing a functional that quantifies the number of “bad” particles through a weighted sum over all possible counts of such particles. This constitutes a first step toward a more general extension in non-exchangeable quantum systems.
Related literature
@@
The earliest works on mean-field limits and the related question of propagation of chaos for exchangeable interacting particle systems go back to Kac [38]. Since then, a vast literature has developed; see [24, 25] or [52] for comprehensive reviews. A growing area of interest is the mean-field limit for non-exchangeable interacting particle systems. Utilizing graphon theory, many works have established mean-field limits for dense graphs (see [27, 4, 18, 9]) and for sparse graphs (see [29, 37]). Within the graphon framework, extensions of classical mean-field games have also been proposed, see [2, 21, 46, 3, 55]. In the quantum setting, the notion of quantum propagation of chaos for exchangeable systems dates back to Spohn [64], and substantial literature has followed, see, for example, [1, 8, 47, 60, 61, 41]. In the static case, [68, 63] investigate the mean-field limit under graphon-type interactions. However, they do not address the question of propagation of chaos. Our work can be seen as a first attempt in that direction.
Organization
@@
The paper is organized as follows. In Section 2, we recall the main notions from quantum filtering theory, graphon theory, and the Fubini extension. Section 3 presents the formalism of non-exchangeable observed interacting quantum particles. In Section 4, we introduce the limiting graphon quantum filtering system in a heuristic manner and then study its well-posedness. Section 5 is devoted to the rigorous derivation of this system for blockwise interactions under perfect measurement efficiency. Section 6 discusses applications to graphon-based quantum dynamic games and quantum state preparation. Section 7 concludes the paper with some perspectives for future research. Appendix A discusses the principle of indistinguishability in quantum mechanics, while Appendix B contains technical results concerning the linear and nonlinear equations.
Notations and facts
@@
The sets of natural, integer, real and complex numbers are denoted $\mathbb{N},\mathbb{Z},\mathbb{R},\mathbb{C}$, respectively. For any $m\in\mathbb{C}$, $\overline{m}$ denotes its complex conjugate and the roman symbol $\mathrm{i}$ is used for pure imaginary number.
@@
Let $T>0$ be a fixed finite time horizon. Given a Polish space $S$, let $\mathcal{C}([0,T],S)$ space of continuous from $[0,T]$ to $S$ equipped with the appropriate topology.
The set of all permutation from $\{1,\dots,N\}\to\{1,\dots,N\}$ is denoted $\mathfrak{S}_{N}$.
In this paper we use $C$ to denote various constants and $C(p)$ to emphasize the dependence on some parameter $p$. Their values may change from line to line.
We fix the notation $\mathbb{H}$ for complex separable Hilbert space, endowed with the scalar product $\langle.,.\rangle:\mathbb{H}\times\mathbb{H}\to\mathbb{C}$, associated with norm $\|\psi\|:=\sqrt{\langle\psi,\psi\rangle}$. We shall use both the notation $\psi$ and the “ket” |$\ket{\psi}$ for a vector in $\mathbb{H}$, and $\bra{\psi}$ | the “bra” for an element of the dual space $\mathbb{H}^{\star}\simeq\mathbb{H}$, thus $\braket{\psi}{\psi^{\prime}}:=\braket{\psi,\psi^{\prime}}$ is the inner product between ${\psi}$ and ${\psi^{\prime}}$.
We define set $\mathcal{B}(\mathbb{H})$ for the space of bounded operators with norm
@@
often we omit the subscript ”op” and just denote $\|O\|$. We denote by ${{O}}^{\dagger}$ the adjoint operator of any operator $O\in\mathcal{B}(\mathbb{H})$, and we denote $\langle O\rangle_{\psi}:=\bra{\psi}O\ket{\psi}$.
For any operator ${O}\in\mathcal{B}(\mathbb{H})$ and for $l\in\{1,\dots,N\}$ denote $\mathbf{O}_{l}:=\mathbf{1}\otimes\dots\otimes{O}\otimes\dots\otimes\mathbf{1}$ the operator on $\mathcal{B}(\mathbb{H}^{\otimes N})$ that acts only on the $l$-th Hilbert space.
For any operator $B\in\mathcal{B}(\mathbb{H}\otimes\mathbb{H})$ and for $l,l^{\prime}\in\{1,\dots,N\}$ denote $\mathbf{B}_{ll^{\prime}}$ the operator on $\mathcal{B}(\mathbb{H}^{\otimes N})$ that acts only on $\mathbb{H}_{l}$ and $\mathbb{H}_{l^{\prime}}$.
We set the notation $\mathcal{B}_{1}(\mathbb{H})$ for trace-class operators and $\mathcal{B}_{2}(\mathbb{H})$ for Hilbert-Schmidt operators.
The Liouville space $\mathcal{L}_{2}(\mathbb{H})$ is the Hilbert space of all Hilbert-Schmidt operators on $\mathbb{H}$ endowed with scalar product
@@
for all $O_{1},O_{2}\in\mathcal{B}_{2}(\mathbb{H})$, where $\{\psi_{n}\}$ is the orthonormal basis of $\mathbb{H}$.
In quantum mechanics, a state is described by an operator in the space of density operators $\mathcal{S}(\mathbb{H})$, which is defined as:
@@
An important special case of a quantum state is what we call a pure state (or wavefunction state), which represents a state of maximal knowledge. A state $\rho\in\mathcal{S}(\mathbb{H})$ is said to be pure if it can be written as a rank-one projector, that is, there exists a vector $\psi\in\mathbb{H}$ such that $\rho=\ket{\psi}\bra{\psi}$ |.

### v3

@@ Abstract
We consider a non-exchangeable system of interacting quantum particles with mean-field type interactions, subject to continuous measurement on dense graphs. In the mean-field limit, we derive a graphon-based quantum filtering system, establish its well-posedness, and prove propagation of chaos for multi-class bosonic systems with blockwise interactions. We then discuss applications to quantum state preparation and quantum graphon games.
Keywords: interacting particle systems; quantum filtering; graphon mean-field limits; stochastic networks; non-exchangeable particles
Mathematics Subject Classification: 81S25, 81S22, 60K35, 60H30, 49N80
@@ 1 Introduction
The study of large systems of interacting quantum particles is fundamental for understanding collective behavior in microscopic and mesoscopic regimes. As the number of particles increases, directly solving the many-body Schrödinger equation becomes infeasible due to the exponential growth of the underlying Hilbert space. Mean-field limits provide an effective alternative by replacing full interacting dynamics with equations that capture averaged interactions [70, 34, 51]. Establishing rigorous connections between exact many-body dynamics and their mean-field approximations is a central problem in mathematical physics. Beyond their conceptual importance, mean-field models increasingly support emerging quantum technologies, guiding the control of large-scale quantum simulators, informing the development of efficient quantum algorithms, and providing theoretical foundations for scalable quantum architectures.
Most existing mean-field results focus on closed quantum systems and deterministic macroscopic observables [76, 62]. In contrast, many physical platforms are open, continuously monitored, and subject to measurement backaction [81, 79], making the development of a mean-field theory at the level of quantum trajectories both conceptually important and technologically motivated. A further source of complexity comes from the interaction structure. While most of the mean-field literature assumes homogeneous or exchangeable interactions, many real systems exhibit heterogeneous coupling patterns [76]. From a mathematical standpoint, this setting is considerably more challenging: the mean-field limit for non-exchangeable quantum systems has not yet been rigorously established, and the limiting description leads to a continuous family of coupled operator-valued stochastic differential equations whose well-posedness requires new probabilistic tools.
In this paper, we develop a formal framework for systems of quantum particles subjected to continuous measurement and their mean-field graph limits. Such models naturally give rise to dynamic quantum games and, by extension, quantum mean-field games (QMFGs) [46, 47], but their relevance extends far beyond game-theoretic applications.
In standard closed quantum mechanics, the dichotomy between the state evolution (Schrödinger equation) and the measurement postulate (Born rule) prevents the observation of significant dynamics, a phenomena known as the Zeno effect [58]. To overcome this, one must consider open quantum systems subjected to indirect measurement, where the natural way to incorporate continuous measurements is provided by the quantum filtering framework [15, 21, 22]. The a posteriori state evolution is then described by a a stochastic modification of the Schrödinger equation, known as the Belavkin or quantum filtering equation. Depending on the nature of the measurement scheme, two types of quantum filtering equations arise:
@@
    $\mathrm{d}\rho_{t}$ $=-\mathrm{i}[\tilde{H},\rho_{t}]\mathrm{d}t+\left(L\rho_{t}L^{\dagger}-\tfrac{1}{2}\{L^{\dagger}L,\rho_{t}\}\right)\mathrm{d}t+\sqrt{\eta}\left(L\rho_{t}+\rho_{t}L^{\dagger}-\mathrm{t}\mathrm{r}\big((L+L^{\dagger})\rho_{t}\big)\rho_{t}\right)\mathrm{d}W_{t},$
    where $(W_{t})_{t\geq 0}$ is a Wiener process (see Section 2.1 for the meaning of the other terms), with the observation process
@@
    where $({N}_{t})_{t\geq 0}$ is a counting process whose stochastic intensity is given by $\int_{0}^{t}\mathrm{t}\mathrm{r}(L\rho_{s}L^{\dagger})\mathrm{d}s$ and the observation process is $Y_{t}=N_{t}.$
The quantum filtering equations can be derived using quantum probability and quantum stochastic calculus [61, 21]. However, they may also be obtained within standard quantum mechanics as suitable continuous-time limits of repeated measurement schemes [65, 63, 64], or through the theory of quantum instruments [7, 9]. In this paper we focus only on the diffusive case.
In his seminal series of papers, Kolokoltsov [45, 47] developed the quantum mean-field games using the $N$-body quantum filtering equation as a starting point:
$\mathrm{d}\boldsymbol{\rho}_{t}^{N}$ $=-\mathrm{i}{\Big[\sum_{l=1}^{N}\mathbf{\tilde{H}}_{l}+\frac{1}{N}\sum_{l=1}^{N}\sum_{l^{\prime}>l}\mathbf{A}_{ll^{\prime}},\boldsymbol{\rho}_{t}^{N}\Big]}\mathrm{d}t+\sum_{l=1}^{N}\Big({\bf L}_{l}\boldsymbol{\rho}_{t}^{N}{\bf L}^{\dagger}_{l}-\frac{1}{2}\big\{{\bf L}^{\dagger}_{l}{\bf L}_{l},\boldsymbol{\rho}_{t}^{N}\big\}\Big)\mathrm{d}t$
@@
where $\boldsymbol{\rho}_{t}^{N}$ denotes the state of the $N$-particle system evolving on a high-dimensional tensor-product Hilbert space. In contrast with the classical setting, due to entanglement between quantum particles, we cannot separate the individual dynamics of the particles, and the concept of an empirical measure becomes meaningless in the quantum setting. Therefore, one must adopt a different methodology to establish propagation of chaos results. Kolokoltsov adapted the Pickl method [66] to the stochastic setting to derive the corresponding limiting equation
$\mathrm{d}\gamma_{t}$ $=-\mathrm{i}[\tilde{H}+A^{\mathbb{E}_{\mathbb{P}}[\gamma_{t}]},\gamma_{t}]\mathrm{d}t+(L\gamma_{t}L^{\dagger}-\tfrac{1}{2}\{L^{\dagger}L,\gamma_{t}\})\mathrm{d}t+\big(L\gamma_{t}+\gamma_{t}L^{\dagger}-\mathrm{tr}\!\big((L+L^{\dagger})\gamma_{t}\big)\gamma_{t}\big)\mathrm{d}W_{t}.$
@@
This leads to a new McKean–Vlasov type equation formulated on an infinite-dimensional complex-valued Hilbert space. The analysis in [45, 47] addresses only homogeneous interactions. The present work develops the corresponding theory for heterogeneous interactions, allowing the coupling strengths between particles to vary.
Primary contributions.
We first develop a mathematical framework for continuously observed, interacting quantum systems with heterogeneous coupling and establish their mean-field graph limits. Using graphon theory, we derive the graphon-based quantum filtering system
@@
$\mathrm{d}\gamma_{t,u}$ $=-\mathrm{i}\Bigl[\tilde{H}+\int_{0}^{1}w(u,v)A^{\mathbb{E}_{\mathbb{P}}[\gamma_{t,v}]}\mathrm{d}v,\gamma_{t,u}\Bigr]\mathrm{d}t+\Bigl(L\gamma_{t,u}L^{\dagger}-\tfrac{1}{2}\{L^{\dagger}L,\gamma_{t,u}\}\Bigr)\mathrm{d}t$
$\quad+\sqrt{\eta}\Bigl(L\gamma_{t,u}+\gamma_{t,u}L^{\dagger}-\mathrm{tr}\bigl((L+L^{\dagger})\gamma_{t,u}\bigr)\gamma_{t,u}\Bigr)\mathrm{d}W^{u}_{t},\quad u\in I:=[0,1],$
where $\mathbf{W}=\{W^{u}:u\in I\}$ is a collection of essentially pairwise independent Wiener processes, and $w$ is a graphon function $w:I\times I\to I$, indicating the strength of the pairwise interaction. Establishing the well-posedness of this system requires addressing some obstacles that do not appear in previous cases dealing with Belavkin diffusive equation.
The first difficulty lies in identifying an appropriate probability space on which to formulate our system of stochastic differential equations. Since the system is driven by a continuum family of independent Wiener processes, serious measurability issues arise. In fact, it is well known in the literature that a continuum family of independent Wiener processes cannot be constructed on a standard probability space while preserving joint measurability (see, for instance, [42] or [25, Section 3.7]). To overcome this difficulty, we specify a probability space that ensures the joint measurability of $\mathbf{W}$ via a Fubini extension [73] which extended the usual product space.
The second difficulty is to identify an appropriate valued space. Indeed, although each $\gamma_{t,u}$ is a trace-class operator, the space of trace-class operators $\mathcal{B}_{1}(\mathbb{H})$ forms a Banach space where Itô theory is ill-defined. Extensions of stochastic calculus via UMD theory [80] (the framework for Banach spaces with unconditional martingale differences) do not apply here, as the space of trace-class operators does not belong to this category. Consequently, following the approach of [48], we work within the larger space of Hilbert–Schmidt operators $\mathcal{B}_{2}(\mathbb{H})$. Equipped with the appropriate inner product, this forms the Liouville space $\mathcal{L}_{2}(\mathbb{H})$, which is a Hilbert space [54]. This allows us to invoke established results for Hilbert space-valued SDEs. The proof then proceeds as follows. Since the trace operation is unbounded on the Liouville space, standard Lipschitz arguments cannot be applied directly. To overcome this, we parametrize the law dependence by a deterministic function and linearize the system to obtain a strong solution via standard arguments for Hilbert space-valued SDEs. We then apply a version of Girsanov transformation on the Fubini extension space to construct a weak solution for parametrized non-linear system. For the delicate issue of positivity, we extend to the infinite-dimensional Hilbert space the argument of [56] which consists in dealing with the complete observed evolution and using conditional expectation to project into the imperfect observed equation. We then prove pathwise uniqueness and invoke the Hilbert space formulation of the Yamada–Watanabe theorem to obtain a strong solution. Finally, to complete the well-posedness result, we establish stability result with respect to the underlying graphon function.
@@
A second major objective of the paper is to establish the trajectory-level mean-field limit for non-exchangeable open quantum systems. To the best of our knowledge, such limits have not been studied in the literature. We consider the block-wise graphon case under perfect measurement efficiency $(\eta=1)$, where the limiting system can be described by a representative trajectory for each block. We extend Pickl’s method by introducing a weighted functional that quantifies the total number of “bad” particles across all blocks. This provides, for the first time, a propagation-of-chaos type result for heterogeneous open quantum systems, providing an initial step toward a general theory of mean-field limits in non-exchangeable quantum settings.
Finally, we discuss applications. We first examine quantum state preparation through continuous measurement and feedback, showing that graphon quantum filtering significantly reduces the complexity of feedback-based stabilization in many-body systems. In particular, each particle’s state converges to an eigenstate of the measurement operator, and appropriate feedback laws stabilize the population toward any desired eigenstate. We then introduce the concept of a graphon quantum dynamic game, demonstrating how the limiting model extends classical mean-field game ideas. The paper concludes with several perspectives and directions for future research.
Related literature.
@@
The study of mean-field limits and the associated notion of propagation of chaos for exchangeable interacting particle systems dates back to the seminal work of Kac [43]. Since then, a vast literature has developed; see [26, 27] or [57] for comprehensive reviews. A growing area of interest is the mean-field limit for non-exchangeable interacting particle systems. Using graphon theory, many works have established mean-field limits for dense graphs (see [29, 6, 20, 11]) and for sparse graphs (see [31, 41]). Within the graphon framework, extensions of classical mean-field games have also been proposed, see [3, 23, 50, 5, 60]. In the quantum setting, the notion of quantum propagation of chaos for exchangeable systems dates back to Spohn [70], and substantial literature has followed, see, e.g., [1, 10, 51, 66, 67, 45]. From a physics perspective and in the static case, [76, 69] examined graphon-type interaction structures in spin closed systems, but without addressing the question of the justification of mean-field limit. The present work can be viewed as a first step in that direction.
Organization.
@@
The paper is organized as follows. In Section 2, we recall the main notions from quantum filtering theory, graphon theory, and the Fubini extension. Section 3 presents the formalism of non-exchangeable observed interacting quantum particles. In Section 4, we introduce the limiting graphon quantum filtering system in a heuristic manner and then study its well-posedness. Section 5 is devoted to the rigorous derivation of this system for blockwise interactions under perfect measurement efficiency. Section 6 discusses applications to quantum state preparation and graphon-based quantum dynamic games. Section 7 concludes the paper with some perspectives for future research. Appendix A discusses the principle of indistinguishability in quantum mechanics, while Appendix B contains technical results concerning the linear and nonlinear equations.
Some notations and facts.
@@
The sets of natural, integer, real, and complex numbers are denoted by $\mathbb{N},\mathbb{Z},\mathbb{R},\mathbb{C}$, respectively. For any $\in\mathbb{C}, z¯\overline{z}$ denotes its complex conjugate, and the roman symbol $\mathrm{i}$ is used for the pure imaginary unit. We write $\mathbb{I}_{A}$ for the indicator of a set $A$. For a probability measure $\nu$, $\mathbb{E}_{\nu}$ denotes expectation under $\nu$. For $k\in\mathbb{N}$, we use the shorthand $k]:=\{1,\dots,k\}$ and denote by $\mathfrak{S}_{k}$ the set of all permutations of $[k]$.
@@
Let $T>0$ be a fixed finite time horizon. For a Banach space $E$, let $\mathfrak{B}(E)$ denote its Borel $\sigma$-field. We define $\mathcal{C}_{E}:=\mathcal{C}([0,T],E)$ as the space of continuous functions from $[0,T]$ to $E$, endowed with the topology of uniform convergence.
For a measure space $(J,\Sigma,\mathrm{r})$, define
$L^{2}_{E}(J,\mathrm{r}):=\Big\{X:J\to E\mid X\;\text{is }\Sigma\text{--}\mathfrak{B}(E)\text{-measurable and }\int_{J}\|X(s)\|_{E}^{2}\mathrm{r}(\mathrm{d}s)<\infty\Big\}.$
As usual, we identify elements of $L^{2}_{E}(J,\mathrm{r})$ that are equal $\mathrm{r}$-almost everywhere (a.e.).
[…]
@@
[…]
The commutator and anti-commutator are defined as follows: for $O_{1},O_{2}\in\mathcal{B}(\mathbb{H})$,
[…]
For any operator $O\in\mathcal{B}(\mathbb{H})$ and $l\in\{1,\dots,N\}$, we denote by
$\mathbf{O}_{l}:=\mathbf{1}\otimes\dots\otimes{O}\otimes\dots\otimes\mathbf{1},$
the operator in $\mathcal{B}(\mathbb{H}^{\otimes N})$ that acts as $O$ on the $l$-th tensor factor and as the identity on all the others (i.e., an operator on the Hilbert space $\mathbb{H}^{\otimes N}$).
For any operator $B\in\mathcal{B}(\mathbb{H}\otimes\mathbb{H})$ and $l,l^{\prime}\in\{1,\dots,N\}$, we denote by $\mathbf{B}_{ll^{\prime}}$ the operator on $\mathcal{B}(\mathbb{H}^{\otimes N})$ that acts only on $\mathbb{H}_{l}$ and $\mathbb{H}_{l^{\prime}}$. We denote by $\mathcal{B}_{1}(\mathbb{H})$ the space of trace-class operators and by $\mathcal{B}_{2}(\mathbb{H})$ the space of Hilbert–Schmidt operators:
$O\in\mathcal{B}_{1}(\mathbb{H})\Longleftrightarrow\mathrm{t}\mathrm{r}(|O|)<\infty,\ O\in\mathcal{B}_{2}(\mathbb{H})\Longleftrightarrow\mathrm{t}\mathrm{r}(OO^{\dagger})<\infty.$
The Liouville space $\mathcal{L}_{2}(\mathbb{H})$ is the Hilbert space of all Hilbert–Schmidt operators on $\mathbb{H}$ endowed with the scalar product
@@
for all $O_{1},O_{2}\in\mathcal{B}_{2}(\mathbb{H})$, where $\{\psi_{n}\}$ is an orthonormal basis of $\mathbb{H}$.
In quantum mechanics, a state is described by an operator in the space of density operators $\mathcal{S}(\mathbb{H})$, defined as
@@
An important special case of a quantum state is a pure state (or wavefunction state), which represents a state of maximal knowledge. A state $\rho\in\mathcal{S}(\mathbb{H})$ is said to be pure if it can be written as a rank-one projector, i.e., there exists a vector $\psi\in\mathbb{H}$ such that $|\rho=\ket{\psi}\bra{\psi}$.
Throughout the paper, we use $C$ to denote various constants and $C(p)$ to emphasize dependence on some parameter $p$. Their values may change from line to line.